One mole of oxygen at 273 K and atmospheric pressure is adiabatically

compressed to 5 atm. Calculate the final temperature. Also calculate the work

done on the gas. Take gama = 1.4 and R = 8.31 J mol-1 K-1.


1
Expert's answer
2021-03-10T17:21:50-0500

Solution:


Adiabatically compression of O₂:  P₁V₁¹·⁴ = P₂V₂¹·⁴


1) Combined gas law: "\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"


2) Then, "\dfrac{P_1^{1.4}V_1^{1.4}}{T_1^{1.4}}=\dfrac{P_2^{1.4}V_2^{1.4}}{T_2^{1.4}}"


"\dfrac{2}{1}: \dfrac{P_1^{0.4}}{T_1^{1.4}}=\dfrac{P_2^{0.4}}{T_2^{1.4}}"


Then, "\sqrt[1.4]{T₁¹·⁴ × (\frac{P₂}{P₁})⁰·⁴}"


Final temperature, "\sqrt[1.4]{273^{1.4}[\frac{5}{1}]^{0.4}}=432 K"


Work done one the gas = ΔPV = nRΔT = 1 × 8.31 × (432 - 273) W = 573 J